QUESTION 30 In the following binomial situation, identify q. Suppose 80% of all new cars come with cars at a dealership, what is the probability that 300 of them come with moon roofs? moon roofs. If there are 320 new OA. q 0.9375 ОВ.q:0.80 OC.q 0.0625 O D.q 0.20 QUESTION 31 Inthe following binomial situation, identify X.-15% of a doctors patients have diabetes. In a sample of 10 new patients, what is the probability that at least 8 of them will have diabetes? O A. x 10 O B. X 0, 1, 2, 3, 4, 5, 6,7,8 O D. X - 8, 9,10
Answer #1

30) here as probability of given event p=0.8

therefore q=1-p=1-0.8 =0.2

option D

31)

here as at least 8 means ; 8 or 9 or 10

therefore X=8,9,10

option D